OpenStudy (anonymous):

If the domain of the function y= -4x+3 is -2< x <= 3 what is the range

OpenStudy (anonymous):

heeeeeelppp

OpenStudy (anonymous):

Is this calculus?

OpenStudy (anonymous):

no, this is algebra II

OpenStudy (anonymous):

Okay, so start with \[ -2<x\leq 3 \]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

What you want to end up with is something like \[ a<-4x+3<b \]

OpenStudy (anonymous):

What does y= -4x -2 has to do with inequalities

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you please explain me how did you mix the inequaility with the equation? I'm on 11th grade xD

OpenStudy (anonymous):

Okay so we start out with \[ -2< x \leq 3 \]Right? And we have to change it so that in the middle is \(y\). This means we need to get \(x\to -4x+3\)

OpenStudy (anonymous):

so it takes two steps: 1) multiply by \(-4\) 2) add \(3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ −2<x≤3 \]Multiply it by \(-4\), this causes inequalities to flip since it is a negative number\[ −2\times -4>-4x \geq 3\times -4 \]Now simplify\[ 8>-4x\geq -12 \] This completes step 1)

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

Next step is add \(3\). Can you do it?